解析的整数論の新しい展開

New Aspects of Analytic Number Theory

京都大学数理解析研究所の共同研究事業の一つとして、下記のように研究集会を催しますので、ご案内申し上げます。なお、この研究集会は森田康夫氏代表の科研費から援助を受けております。

研究代表者谷川好男 (名大 · 多元数理）

記

日時： 2001 年 11 月 26 日（月） 10:00 〜 11 月 30 日（金） 16:30

場所：京都大学数理解析研究所４階 420 号室 京都市左京区北白川追分町

プログラム

11 月 26 日（月）

10:00〜10:50 松本耕二 (名大 · 多元数理)

二重ガンマ関数とある種の実二次体

11:00〜11:50 K. Srinivasa Rao (Institute of Mathematical Sciences, India）

Quantum theory of angular momentum, hypergeometric

series and polynomial zeros

13:30〜14:00 藤森雅巳（神奈川工科大 · 工）

線形不等式系の族について

14:10〜14:50 大野泰生（近畿大 · 理工）
      多重ゼータ値の線形関係式について
15:20〜15:50 佐藤晋 (東大 · 数理）
      Beukers 積分とその一般化について
16:00〜16:30 市原由美子 (名大 · 多元数理)

Rankin-Selberg L 関数における Siegel-Tatuzawa の定理について

16:40〜17:20 知念宏司 (大阪工業大 · 工), 村田玲音 (明治学院大 · 経済)

a (mod p) の剰余位数の分布について, II 11 月 27 日（火）

9:30〜10:20  野口潤次郎 (東大 · 数理）
     Nevanlinna Theory と abc-Conjecture
10:30〜11:20 Carl Pomerance (Bell Laboratories, USA）
      Primitive roots
11:40〜12:20 平田典子（日大 · 理工）
      ディオファントス連立不等式におけるギャップ原理
14:00〜14:30 永田誠 (京大 · 数理研)
      G-関数に関するディオファンタス近似について
14:40〜15:20 小松尚夫 (三重大 · 教育)
      新しい規則性を示す連分数展開 – Tasoev の連分数
15:40〜16:10 橋本竜太 (名大 · 人間情報)
      不定方程式 X^{3} − (c^{3} − 1)Y^{3} = N と^{3} \sqrt{c^{3} − 1} の連分数展開
16:20〜16:50 鹿子智明（慶応大 · 理工）
      線形回帰数列のある種の逆数和の超越性
17:00〜17:30 名越弘文 (京大 · 数理研)

保型 L-関数の臨界値と中心極限定理

11 月 28 日（水）

9:30〜10:20 Bruce C. Berndt (Univ. of Illinois, USA)
      Ramanujan’s work on Eisenstein series, especially

in his lost notebook

10:30〜11:20 本橋洋一（日大 · 理工）
      表現論における一般ベッセル函数と解析的整数論の諸問題
11:30〜12:20 Georges Rhin (Univ. de Metz, France）
      On irrationality measures of values of the zeta function

at rational integers

午後自由時間

11 月 29 日 (木)

9:30〜10:20 Kalman Györy (Univ. of Debrecen, Hungary）
      On the number of solutions of decomposable

polynomial equations

10:30〜11:20 北岡良之（名城大 · 理工）
      代数体の単数の分布について
11:40〜12:20 秋山茂樹 (新潟大 · 理)

標準数系の分類に関する最近の進展

14:00〜14:30 小松和志 (高知大 · 理)

射影法による非周期タイル貼りについて

14:40〜15:10 天羽雅昭 (群馬大 · 工)

q-超幾何級数の特殊値の１次独立性

15:30〜16:00 水野義紀 (阪大 · 理学)

コーエン型アイゼンシュタイン級数の Rankin-Selberg convolution

16:10〜16:40 Temenoujka Peneva (筑波大 · 数学)

An additive problem with Piatetski-Shapiro primes and almost-primes

16:50〜17:30 三河寛 (筑波大 · 数学)

素数によるヂィオファントス近似について

レセプション 18:30〜 於 芝蘭会館

11 月 30 日（金）

9:30〜10:20 Tarlok N. Shorey (Tata Institute Fundamental Research, India)
     Powers in arithmetic progressions
10:30〜11:20 Winfried Kohnen (Universität Heidelberg, Germany)
      On special values of the j-function and related modular functions
11:40〜12:20 江上繁樹 (富山大 · 工)
      Diophantine Dirichlet series

14:00〜14:30 川田浩一 (岩手大 · 教育）
      素数を変数とするいくつかの加法的問題について
14:40〜15:10 Zhai Wenguang (山東師範大-名大）
      On the simultaneous distribution of the fractional parts

of different powers of prime numbers

15:20〜15:50 吉元昌巳 (京大 · 数理研)

マデルング定数とエプシュタインゼータ関数

16:00〜16:30 神谷諭一 (名大 · 多元数理)
      ある Dirichlet 級数の族に関する近似について

New Aspects of Analytic Number Theory

Date ： 26, November – 30, November (2001)

Place : Research Institute of Mathematical Science (RIMS) Kyoto Univ. 4F, Room 420

Organizer : Yoshio Tanigawa (Nagoya Univ.)

PROGRAM

Monday 26, Nov.

10:00〜10:50 Kohji Matsumoto (Nagoya Univ.)

Double gamma functions and certain real quadratic fields

11:00〜11:50 K. Srinivasa Rao (Institute of Mathematical Sciences, India）

Quantum theory of angular momentum, hypergeometric series and polynomial zeros

13:30〜14:00 Masami Fujimori (Kanagawa Inst. of Tech.)

On systems of linear inequalities

14:10〜14:50 Yasuo Ohno (Kinki Univ.)
      On linear relations among multiple zeta values
15:20〜15:50 Susumu Sato (Univ. of Tokyo）
      On Beukers integrals and their generalization
16:00〜16:30 Yumiko Ichihara (Nagoya Univ.)

An analogue of the Siegel-Tatuzawa theorem for Rankin-Selberg L-functions

16:40〜17:20 Koji Chinen (Osaka Inst. of Tech.), Leo Murata (Meiji Gakuin Univ.)

On a distribution property of the residual order of a (mod p), II Tuesday 27, Nov.

9:30〜10:20  Junjiro Noguchi (Univ. of Tokyo）
     Nevanlinna Theory and abc-Conjecture
10:30〜11:20 Carl Pomerance (Bell Laboratories, USA）
      Primitive roots
11:40〜12:20 Noriko Hirata （Nihon Univ.）
      Gap principle in Diophantine systems
14:00〜14:30 Makoto Nagata (Kyoto Univ.)
      On Diophantine approximations related to G-functions
14:40〜15:20 Takao Komatsu (Mie Univ.)
      Continued fraction expansions showing new regularity– Tasoev’s

continued fractions

15:40〜16:10 Ryuuta Hashimoto (Nagoya Univ.)
      Diophantine equation X^{3} − (c^{3} − 1)Y^{3} = N and continued fraction

of ^{3} \sqrt{c^{3} − 1}

16:20〜16:50 Tomoaki Kanoko (Keio Univ.）
      Transcendence of certain reciprocal sums of linear recurrences
17:00〜17:30 Hirofumi Nagoshi (Kyoto Univ.)

Critical values of automorphic L-functions and a central limit theorem

Wednesday 28, Nov.

9:30〜10:20 Bruce C. Berndt (Univ. of Illinois, USA)
      Ramanujan’s work on Eisenstein series, especially in his lost notebook
10:30〜11:20 Yoichi Motohashi （Nihon Univ.）
      Generalized Bessel Functions in the Representation Theory and Problems

in Analytic Number Theory

11:30〜12:20 Georges Rhin (Universite de Metz, France）
      On irrationality measures of values of the zeta function at rational integers
Afternoon  Free time

Thursday 29, Nov.

9:30〜10:20 Kalman Györy (Univ. of Debrecen, Hungary）
      On the number of solutions of decomposable polynomial

equations

10:30〜11:20 Yoshiyuki Kitaoka （Meijyou Univ.）
      Distribution of units of algebraic number fields
11:40〜12:20 Shigeki Akiyama (Niigata Univ.)

Recent progress on classification of canonical number systems

14:00〜14:30 Kazushi Komatsu (Kouchi Univ.)

On a periodic tilings by the projection method

14:40〜15:10 Masaaki Amou (Gunma Univ.)

Linear independence of the values of q-hypergeometric series

15:30〜16:00 Yoshinori Mizuno (Osaka Univ.)

Rankin-Selberg convolution for Cohen type Eisenstein series

16:10〜16:40 Temenoujka Peneva (Tsukuba Univ.)

An additive problem with Piatetski-Shapiro primes and almost-primes

16:50〜17:30 Hiroshi Mikawa (Tsukuba Univ.)

Small values of αp (mod 1)

Reception at Shiran Kaikan 18:30〜 Friday 30, Nov.

9:30〜10:20 Tarlok N. Shorey (Tata Institute Fundamental Research, India)
     Powers in arithmetic progressions
10:30〜11:20 Winfried Kohnen (Universität Heidelberg, Germany)
      On special values of the j-function and related modular functions
11:40〜12:20 Shigeki Egami (Toyama Univ.)
      Diophantine Dirichlet series

14:00〜14:30 Koichi Kawada (Iwate Univ.）
      On ternary additive problems in prime number theory,

and related problems

14:40〜15:10 Zhai Wenguang （Shandong Teachers’ Univ. -Nagoya Univ.）
      On the simultaneous distrubution of the fractional parts

of different powers of prime numbers

15:20〜15:50 Masami Yoshimoto (Kyoto Univ.)

Madelung constants and Epstein zeta functions

16:00〜16:30 Yuichi Kamiya (Nagoya Univ.)
      An approximation for certain Dirichlet sereis